Find the area bounded by the curve x=y²+2y and the line x=3.

1
Expert's answer
2022-05-26T17:20:17-0400

Solution

Points of intersection of the given curves are solution of equation y2 + 2y = 3   =>  y2 + 2y – 3 = 0 => y1 = -3, y2 = 1.

So area between curves is

"A=\int_{-3}^{1}\left[3\ -y^2-2y\right]dy=\left[3y\ -\frac{1}{3}y^3-y^2\right]\left|\begin{matrix}1\\-3\\\end{matrix}\right."

A = 3 – 1/3 – 1 + 9 – 9 + 9 = 11 – 1/3 = 10.66667